# Morphology-Consistent Humanoid Interaction through Robot-Centric Video Synthesis

### 20 Mar 2026 v1

@@ Abstract
Equipping humanoid robots with versatile interaction skills typically requires either extensive task-specific policy training or explicit human-to-robot motion retargeting. However, learning-based policies are hindered by prohibitive data collection costs, limiting their scalability. Meanwhile, retargeting paradigms rely heavily on human-centric pose estimation (e.g., SMPL), which inevitably introduces the morphology gap. Such skeletal scale mismatches result in severe spatial misalignments when mapped to robots, compromising interaction success. In this work, we propose Dream2Act, a robot-centric framework that enables zero-shot interaction through generative video synthesis. Given an image of the predefined robot and target object in third-person view, our framework leverages video generation models to envision videos in which the physical robot completes the task with spatially aligned, morphology-consistent motion. We then employ a robust, high-fidelity pose extraction system to recover physically feasible, robot-native joint trajectories from these synthesized “dreams”, which are subsequently executed via a general-purpose whole-body controller. By operating strictly within the robot-native coordinate space, Dream2Act effectively avoids traditional retargeting errors and eliminates the need for task-specific policy training. We evaluate Dream2Act on the Unitree G1 across four categories of whole-body mobile interaction tasks (multi-position ball kicking, sofa sitting, bag punching, and box hugging). Dream2Act achieves an overall task success rate of $37.5\%$, compared to $0\%$for conventional retargeting pipelines. Retargeting frequently fails to establish correct physical contacts due to the morphology gap, with errors further compounded under locomotion. In contrast, Dream2Act maintains robot-consistent spatial alignment throughout execution, enabling reliable contact formation and substantially higher task completion.
@@ 1 Introduction
Figure 1: Dream2Act: Zero-shot Humanoid Interaction via Robot-Centric Video Synthesis. (A) Traditional human-centric retargeting fails due to the morphology gap, causing spatial misalignments. (B) Our Dream2Act pipeline leverages Seedance 2.0 to hallucinate morphology-consistent interactions directly in the robot’s native space. (C) This visual prior translates into successful zero-shot physical execution for spatially-sensitive tasks like sitting on a sofa and kicking a ball.
The ultimate goal of embodied artificial intelligence is to deploy humanoid robots capable of executing general-purpose interaction tasks in unstructured human environments. To achieve this, acquiring dexterous interaction skills from demonstrations has emerged as a promising paradigm [9, 12, 37]. However, capturing high-quality humanoid demonstrations conventionally relies on either expensive hardware-intensive teleoperation [10] or precise whole-body pose estimation for human-to-robot mapping [25]. Both pipelines are bottlenecked by high hardware costs, substantial human labor, and a lack of scalability.
@@
To exploit human videos, current mainstream approaches heavily rely on human-to-robot motion retargeting [19, 1]. These paradigms typically utilize human-centric kinematic models, such as SMPL [18], to extract poses from human videos and subsequently map them onto the robot-specific embodiment.
Unfortunately, this human-centric pipeline is intrinsically hindered by the morphology gap: humans and humanoids differ in skeletal proportions, degrees of freedom, and joint limits. Forcing human kinematics onto a robotic skeleton inevitably introduces severe spatial misalignments, a problem further exacerbated by the compounding errors of mobility. Even in fundamental whole-body mobile interactions, these scale mismatches and accumulated locomotion drifts severely warp the absolute spatial relationship between the agent and the target, consistently compromising task success.
@@
Subsequently, we employ a high-fidelity pose estimation and reconstruction pipeline to recover robot-native kinematic trajectories directly from these synthesized sequences. Finally, these trajectories are executed through a general-purpose whole-body controller SONIC [20]. By entirely bypassing the human embodiment intermediate, Dream2Act avoids retargeting-induced errors and sidesteps the morphology gap.
Crucially, our framework enables context-aware, task-level zero-shot interaction. By conditioning the generative process on a single real-world observation and a semantic task prompt, Dream2Act achieves the ability to execute novel tasks in a training-free manner, bypassing the need for any prior exposure to specific task dynamics or environment-specific data collection. We comprehensively validate Dream2Act across four representative categories of whole-body mobile interaction tasks deployed on the Unitree G1 humanoid. Extensive physical experiments demonstrate that while conventional optimization-based retargeting baselines suffer catastrophic failure in establishing correct physical contacts, our approach reliably achieves functional spatial alignment and successful physical engagement.
@@
  • We introduce a robust, decoupled kinematic recovery pipeline that effectively bridges the pixel-to-physics gap. By cascading native 2D pose estimation, a domain-specific 3D lifting network, and URDF-constrained inverse kinematics, our approach successfully translates generative visual priors into physically executable joint trajectories.
@@ 2.1 Human-Centric Motion Acquisition and Retargeting
Imitating human demonstrations is a cornerstone for acquiring humanoid robotic skills. A prevalent paradigm relies on a multi-stage pipeline: first, human kinematics are extracted from monocular videos or motion capture data using human-centric parametric models such as SMPL [18] or GVHMR [29]. Similarly, in the Text-to-Motion domain, leading generative models, such as MDM [31], HY-Motion [33], Dart [11] and ViMogen [16] from the T2M benchmark [36], synthesize high-fidelity human kinematics directly from semantic prompts.
Regardless of the source, these human-centric poses must be mapped to robotic embodiments through optimization-based solvers like GMT [7] or GMR [1] to bridge skeletal discrepancies. These retargeted trajectories then serve as reference motions for training RL policies or whole-body trackers, a methodology pioneered by DeepMimic [25] and advanced by recent dexterous systems such as PHC [19],MimicKit [26] OmniH2O [12], and HumanPlus [9].
While effective for general locomotion, these methodologies are fundamentally bottlenecked by the morphology gap. Beyond simple discrepancies in skeletal proportions, these retargeting processes are often optimizing for kinematic resemblance in isolation from the physical scene. This leads to severe spatial misalignments and lost contact precision in object-centric tasks.
@@ 2.3 Robotic Perception and General-Purpose Tracking
Translating generative visual hallucinations into physical robot actions requires a robust integration of robotic pose perception and whole-body tracking [25]. While human pose models like HRNet [30] and ViTPose [35] are mature on datasets like COCO [17], they struggle with robotic embodiments due to morphology gaps and domain-specific data scarcity [35]. Dream2Act overcomes this by leveraging simulation environments like IsaacLab [22] and motion priors such as AMASS [21] to synthesize a native dataset, enabling precise pose perception grounded in the robot’s specific kinematic skeleton.
On the execution side, the field has evolved from task-specific imitation toward general-purpose whole-body tracking. Traditional frameworks often relied on analytical solvers like Pinocchio [6] or specialized systems like TWIST [37]. While agile trackers such as BeyondMimic [15] improved responsiveness, they remain limited to specific motion distributions. Recently, the development of general tracking foundations—such as BFM-Zero [14] and Sonic [20]—has enabled the tracking of diverse, complex motions across unstructured environments. Dream2Act synergizes with these foundations by providing morphology-consistent and kinematically plausible visual priors, achieving stable zero-shot physical execution without requiring additional task-specific training.
@@ 3 Method
We specifically select this model over other alternatives due to its state-of-the-art performance in two critical dimensions required for robotic perception: (1) Physical Plausibility: It exhibits superior adherence to real-world physics (e.g., rigid body dynamics and contact constraints) during complex interactions; and (2) Morphological Consistency: It maintains exceptional structural rigidity of the subjects without introducing catastrophic distortions or artifacts, even under significant viewpoint shifts or prolonged task execution.
Before executing any novel interaction task, our framework only requires a single initial frame captured by a third-person camera in the real environment, containing the physical Unitree G1 robot and the target object. This frame serves as the exact spatial conditioning. By incorporating text instructions or semantic task prompts, Seedance 2.0 synthesizes a physically intuitive and temporally coherent video that shows the robot successfully completing the intended task. Crucially, the generated video effectively maintains a fixed third-person camera perspective and preserves the physical dimensions of the G1 with high morphological fidelity. This high-quality, morphology-consistent video stream thereby provides a reliable visual prior for the subsequent native kinematic recovery.
@@ 3.2 Domain-Specific Robotic Dataset Construction
Simulation-based High-Fidelity Rendering. To construct a robust motion prior, we leverage the AMASS dataset [21]. We first retarget its extensive human motion sequences to the G1 29-DoF kinematic model, yielding a massive initial trajectory collection of 5,337,514 frames. Since highly correlated adjacent frames offer marginal information gain and inherently degrade the learning efficiency of subsequent networks, we perform a rigorous temporal downsampling. This step effectively distills the dataset into 5,338 diverse and representative keyframes for high-fidelity rendering. With the robot anchored at the origin, we construct a hemispherical, omnidirectional multi-camera system in Isaac Lab. This array comprises nine high-resolution virtual cameras: four orthogonal side views, four diagonal views with a 45 downward pitch, and one strict top-down view. To bridge the coordinate system gap between the physics simulation (USD convention) and standard computer vision pipelines, we perform rigorous orthogonal frame transformations. Using the derived camera extrinsics and intrinsics, we efficiently project the 3D world joints into local camera space and subsequently into precise 2D pixel coordinates $u,v]^{T}$. Beyond skeletal joints, we synchronously render depth maps and instance segmentation masks to automatically extract tight bounding boxes and center points for the robot. This pipeline furnishes a rich, multi-modal ground-truth suite for the subsequent visual generation and lifting modules.
In-the-Wild Data and Human-in-the-Loop 2D Annotation. To bridge the Sim-to-Real gap for visual perception, we collect in-the-wild videos of the physical G1 robot and temporally downsample them by a factor of $1/10$ to reduce redundancy. We then partition these processed frames into a training set and an evaluation set. For the training set, we adopt a Human-in-the-Loop (HITL) annotation pipeline: a purely simulation-trained model first pre-annotates the 2D pixel coordinates, followed by expert manual correction of any misaligned joints. Conversely, the evaluation set undergoes multi-round, fine-grained expert 2D annotation from scratch to establish a high-confidence benchmark. Crucially, this real-world dataset is exclusively utilized to fine-tune the 2D pose estimation network (detailed in Sec. 3.3) against real-world illumination and textures, whereas our 2D-to-3D lifting network relies entirely on the exact 3D geometric ground truth provided by the simulation environment.
@@ 3.3 Native Pose Estimation and 2D-to-3D Lifting
Equipped with our hybrid dataset, we employ ViTPose [35] as the backbone architecture for Parameter-Efficient Fine-Tuning (PEFT). Specifically, we freeze the first nine layers of the pre-trained ViT-Base model and fine-tune only the final three layers alongside the top-down Gaussian heatmap head. Given an input image $I$, this estimator accurately predicts the 2D keypoints $P_{2D}$ for the $N=30$ native robotic joints.
@@ 3.4 Morphology-Aware Kinematic Recovery and Physical Execution
To recover the authentic motion trajectory of the G1 robot from the inferred 3D visual keypoints $P_{3D}$, we design a decoupled pipeline that separates internal joint kinematics from global root dynamics, culminating in authentic physical deployment.
@@
subject to joint limits $q_{min}\leq q\leq q_{max}$. By fixing the root node at the origin and optimizing solely on position constraints, the IK solver generates a structurally flawless internal joint configuration $q^{*}$ that strictly adheres to the rigid skeleton of the G1.
Stage 2: Scale Recovery and Root Pose Optimization. We apply Forward Kinematics (FK) to the internal state $q^{*}$ (with the root fixed as an identity transformation) to obtain precise local 3D points $P_{local}\in\mathbb{R}^{N\times 3}$. Subsequently, we align these points with the unnormalized original 2D detections $P_{2D}$ using the Perspective-n-Point (PnP) algorithm:
@@ 4 Experiments
Our experiments are systematically designed to answer the following research questions: (1) Can Dream2Act enable zero-shot physical interactions compared to human-centric retargeting? (2) Does robot-centric video synthesis offer a more physically plausible paradigm for free-space motion generation than conventional text-to-motion models? (3) How do our specific system designs—namely, the hybrid dataset contribute to robust perception?
@@ 4.1 Experimental Setup and Datasets
Datasets Construction. Training highly accurate native pose estimators requires domain-specific data. We construct a hybrid dataset comprising two distinct domains: (1) Simulation Dataset: We leverage AMASS motion priors, retargeting them to the G1 to yield over 5.3 million frames, which are then rigorously downsampled to 5,338 diverse keyframes. Through a 9-camera hemispherical setup in Isaac Lab, we render high-fidelity images accompanied by exact 3D geometric ground truth ($P_{world}$), bounding boxes, and masks. (2) In-the-Wild Real Dataset: To address the Sim-to-Real visual domain shift, we collect real-world videos of the G1, temporally downsampled to a factor of 1/10. These frames are partitioned into a 90% training set (refined via Human-In-The-Loop expert correction) and a 10% evaluation set (341 frames, annotated from scratch via multi-round expert labeling) to serve as a high-confidence benchmark.
Figure 3: G1 Domain-Specific Hybrid Dataset Construction Pipeline. Our approach synergizes cost-free, high-fidelity exact 3D/2D geometry rendered from a 9-camera simulation environment (Isaac Lab) with Real-World Human-in-the-Loop (HITL) annotations to correct visual domain shifts, enabling robust native pose estimation.

### 24 Mar 2026 v2

@@ Abstract
Equipping humanoid robots with versatile interaction skills typically requires either extensive task-specific policy training or explicit human-to-robot motion retargeting. However, learning-based policies are hindered by prohibitive data collection costs, limiting their scalability. Meanwhile, retargeting paradigms rely heavily on human-centric pose estimation (e.g., SMPL), which inevitably introduces the morphological gap. Such skeletal scale mismatches result in severe spatial misalignments when mapped to robots, compromising interaction success. In this work, we propose Dream2Act, a robot-centric framework that enables zero-shot interaction through generative video synthesis. Given an image of the target robot and target object in third-person view, our framework leverages video generation models to synthesize video sequences in which the physical robot completes the task with spatially aligned, morphology-consistent motion. We then employ a robust, high-fidelity pose extraction system to recover physically feasible, robot-native joint trajectories from these synthesized “dreams”, which are subsequently executed via a general-purpose whole-body controller. By operating strictly within the robot-native coordinate space, Dream2Act effectively avoids traditional retargeting errors and eliminates the need for task-specific policy training. We evaluate Dream2Act on the Unitree G1 across four categories of whole-body mobile interaction tasks (multi-position ball kicking, sofa sitting, bag punching, and box hugging). Dream2Act achieves an overall task success rate of $37.5\%$, compared to $0\%$ for conventional retargeting pipelines. Retargeting frequently fails to establish correct physical contacts due to the morphological gap, with errors further compounded under locomotion. In contrast, Dream2Act maintains robot-consistent spatial alignment throughout execution, enabling reliable contact formation and substantially higher task completion.
@@ 1 Introduction
Figure 1: Dream2Act: Zero-shot Humanoid Interaction via Robot-Centric Video Synthesis. (A) Traditional human-centric retargeting fails due to the morphology gap, causing spatial misalignments. (B) Our Dream2Act pipeline leverages Seedance 2.0 to synthesize morphology-consistent interactions directly in the robot’s native space. (C) This visual prior translates into successful zero-shot physical execution for spatially-sensitive tasks like sitting on a sofa and kicking a ball.
The ultimate goal of embodied artificial intelligence is to deploy humanoid robots capable of executing general-purpose interaction tasks in unstructured human environments. To achieve this, acquiring dexterous interaction skills from demonstrations has emerged as a promising paradigm [9, 12, 37]. However, capturing high-quality humanoid demonstrations conventionally relies on either expensive hardware-intensive teleoperation [10] or precise whole-body pose estimation for human-to-robot mapping [25]. Both pipelines are constrained by high hardware costs, substantial human labor, and a lack of scalability.
@@
To exploit human videos, mainstream approaches rely on human-to-robot motion retargeting [19, 1]. These paradigms typically utilize human-centric kinematic models, such as SMPL [18], to extract poses from human videos and map them onto the robot-specific embodiment.
This human-centric pipeline is hindered by the morphology gap: humans and humanoids differ in skeletal proportions, degrees of freedom, and joint limits. Forcing human kinematics onto a robotic skeleton inevitably introduces severe spatial misalignments, a problem compounded by accumulated errors during locomotion. Even in fundamental whole-body mobile interactions, these scale mismatches and accumulated locomotion drifts severely warp the absolute spatial relationship between the agent and the target, consistently compromising task success.
@@
Subsequently, we employ a high-fidelity pose estimation and reconstruction pipeline to recover robot-native kinematic trajectories directly from these synthesized sequences. Finally, these trajectories are executed through a general-purpose whole-body controller SONIC [20]. By entirely bypassing the the intermediate human embodiment representation, Dream2Act avoids retargeting-induced errors and sidesteps the morphology gap.
Our framework enables context-aware, task-level zero-shot interaction. By conditioning the generative process on a single real-world observation and a semantic task prompt, Dream2Act achieves the ability to execute novel tasks in a training-free manner, bypassing the need for any prior exposure to specific task dynamics or environment-specific data collection. We comprehensively validate Dream2Act across four representative categories of whole-body mobile interaction tasks deployed on the Unitree G1 humanoid. Extensive physical experiments demonstrate that while conventional optimization-based retargeting baselines suffer catastrophic failure in establishing correct physical contacts, our approach reliably achieves functional spatial alignment and successful physical engagement.
@@
  • We introduce a robust kinematic recovery pipeline that effectively bridges the pixel-to-physics gap. By cascading native 2D pose estimation, a domain-specific 3D lifting network, and URDF-constrained inverse kinematics, our approach successfully translates generative visual priors into physically executable joint trajectories.
@@ 2.1 Human-Centric Motion Acquisition and Retargeting
Imitating human demonstrations is a cornerstone for acquiring humanoid robotic skills. A prevalent paradigm relies on a multi-stage pipeline: first, human kinematics are extracted from monocular videos or motion capture data using human-centric parametric models such as SMPL [18] or GVHMR [29]. Similarly, in the Text-to-Motion domain, leading generative models, such as MDM [31], HY-Motion [33], Dart [11], and ViMogen [16] from the T2M benchmark [36], synthesize high-fidelity human kinematics directly from semantic prompts.
Irrespective of the source, these human-centric poses must be mapped to robotic embodiments through optimization-based solvers like GMT [7] or GMR [1] to bridge inherent skeletal discrepancies. These retargeted trajectories serve as foundational reference motions for training RL policies or whole-body trackers, a methodology pioneered by DeepMimic [25] and subsequently advanced by dexterous systems such as PHC [19], MimicKit [26] OmniH2O [12], and HumanPlus [9].
While effective for general locomotion, these methodologies are fundamentally bottlenecked by the morphology gap. Beyond simple discrepancies in skeletal proportions, these retargeting processes are often designed to optimize for kinematic resemblance in isolation from the physical scene. This leads to severe spatial misalignments and lost contact precision in object-centric tasks.
@@ 2.3 Robotic Perception and General-Purpose Tracking
Translating generative visual hallucinations into physical robot actions requires a robust integration of robotic pose perception and whole-body tracking [25]. While human pose models like HRNet [30] and ViTPose [35] are mature on datasets like COCO [17], they struggle with robotic embodiments due to the morphology gap and domain-specific data scarcity [35]. Dream2Act overcomes this by leveraging simulation environments like IsaacLab [22] and motion priors such as AMASS [21] to synthesize a native dataset, enabling precise pose perception grounded in the robot’s specific kinematic skeleton.
On the execution side, the field has evolved from task-specific imitation toward general-purpose whole-body tracking. Traditional frameworks often relied on analytical solvers like Pinocchio [6] or specialized systems like TWIST [37]. While agile trackers such as BeyondMimic [15] have improved responsiveness, they remain limited to specific motion distributions. Recently, the development of general tracking foundations—such as BFM-Zero [14] and Sonic [20]—has enabled the tracking of diverse, complex motions across unstructured environments. Dream2Act synergizes with these foundations by providing morphology-consistent and kinematically plausible visual priors, achieving stable zero-shot physical execution without requiring additional task-specific training.
@@ 3 Method
We select this model over alternatives due to its state-of-the-art performance in two critical dimensions required for robotic perception: (1) Physical Plausibility: It exhibits superior adherence to real-world physics (e.g., rigid body dynamics and contact constraints) during complex interactions; and (2) Morphological Consistency: It maintains exceptional structural rigidity of the subjects without introducing distortions or artifacts, even under significant viewpoint shifts or prolonged task execution.
Before executing any novel interaction task, our framework only requires a single initial frame captured by a third-person camera in the real environment, containing the physical Unitree G1 robot and the target object. This frame serves as the exact spatial conditioning. By incorporating text instructions or semantic task prompts, Seedance 2.0 synthesizes a physically intuitive and temporally coherent video that shows the robot successfully completing the intended task. Crucially, the generated video effectively maintains a fixed third-person camera perspective and preserves the physical dimensions of the G1 with high morphological fidelity. This high-quality, morphology-consistent video stream thereby serves as a reliable visual prior for the subsequent native kinematic recovery.
@@ 3.2 Domain-Specific Robotic Dataset Construction
Simulation-based High-Fidelity Rendering. To construct a robust motion prior, we leverage the AMASS dataset [21]. We first retarget its extensive human motion sequences to the G1 29-DoF kinematic model, yielding a massive initial trajectory collection of 5,337,514 frames. Since highly correlated adjacent frames offer marginal information gain and degrade the learning efficiency of subsequent networks, we perform a rigorous temporal downsampling. This step aggressively distills the dataset into 5,338 diverse and representative keyframes for high-fidelity rendering. With the robot anchored at the origin, we construct a hemispherical, omnidirectional multi-camera system in Isaac Lab. This array comprises nine high-resolution virtual cameras: four orthogonal side views, four diagonal views with a 45 downward pitch, and one strict top-down view. To bridge the coordinate system gap between the physics simulation (USD convention) and standard computer vision pipelines, we perform rigorous orthogonal frame transformations. Using the derived camera extrinsics and intrinsics, we efficiently project the 3D world joints into local camera space and subsequently into precise 2D pixel coordinates $u,v]^{T}$. Beyond skeletal joints, we synchronously render depth maps and instance segmentation masks to automatically extract tight bounding boxes and center points for the robot. This pipeline furnishes a rich, multi-modal ground-truth suite for the subsequent visual generation and lifting modules.
In-the-Wild Data and Human-in-the-Loop 2D Annotation. To bridge the Sim-to-Real gap for visual perception, we collect in-the-wild videos of the physical G1 robot and temporally downsample them by a factor of $10$ to reduce redundancy. We then partition these processed frames into a training set and an evaluation set. For the training set, we adopt a Human-in-the-Loop (HITL) annotation pipeline: a purely simulation-trained model first pre-annotates the 2D pixel coordinates, followed by expert manual correction of misaligned joints. Conversely, the evaluation set undergoes multi-round, fine-grained expert 2D annotation from scratch to establish a high-confidence benchmark. Crucially, this real-world dataset is exclusively utilized to fine-tune the 2D pose estimation network (detailed in Sec. 3.3) against real-world illumination and textures, whereas our 2D-to-3D lifting network relies on the exact 3D geometric ground truth provided by the simulation environment.
@@ 3.3 Native Pose Estimation and 2D-to-3D Lifting
Equipped with our hybrid dataset, we employ ViTPose [35] as the backbone architecture for Parameter-Efficient Fine-Tuning (PEFT). We freeze the first nine layers of the pre-trained ViT-Base model and fine-tune only the final three layers alongside the top-down Gaussian heatmap head. Given an input image $I$, this estimator accurately predicts the 2D keypoints $P_{2D}$ for the $N=30$ native robotic joints.
@@ 3.4 Morphology-Aware Kinematic Recovery and Physical Execution
To recover the authentic motion trajectory of the G1 robot from the inferred 3D visual keypoints $P_{3D}$, we design a robust, decoupled pipeline that explicitly separates low-level internal joint kinematics from global root dynamics, culminating in authentic physical deployment.
@@
subject to joint limits $q_{min}\leq q\leq q_{max}$. By fixing the root node at the origin and optimizing solely on position constraints, the IK solver generates a structurally consistent internal joint configuration $q^{*}$ that strictly adheres to the rigid URDF model of the G1.
Stage 2: Scale Recovery and Root Pose Optimization. We apply Forward Kinematics (FK) to the internal state $q^{*}$ (with the root fixed as an identity transformation) to obtain precise local 3D points $P_{local}\in\mathbb{R}^{N\times 3}$. Subsequently, we align these points with the un-normalized original 2D detections $P_{2D}$ using the Perspective-n-Point (PnP) algorithm:
@@ 4 Experiments
Our experiments are systematically designed to answer the following research questions: (1) Can Dream2Act enable zero-shot physical interactions compared to human-centric retargeting? (2) Does robot-centric video synthesis offer a more physically plausible paradigm for free-space motion generation than conventional text-to-motion models? (3) how our specific system designs, namely the hybrid dataset, contribute to robust perception.?
@@ 4.1 Experimental Setup and Datasets
Datasets Construction. Training accurate native pose estimators requires domain-specific data. We construct a hybrid dataset comprising two distinct domains: (1) Simulation Dataset: We leverage AMASS motion priors, retargeting them to the G1 to yield over 5.3 million frames, which are then rigorously downsampled to 5,338 diverse keyframes. Through a 9-camera hemispherical setup in Isaac Lab, we render high-fidelity images accompanied by exact 3D geometric ground truth ($P_{world}$), bounding boxes, and masks. (2) In-the-Wild Real Dataset: To address the Sim-to-Real visual domain shift, we collect real-world videos of the G1, temporally downsampled to a factor of 10. These frames are partitioned into a 90% training set (refined via Human-In-The-Loop expert correction) and a 10% evaluation set (341 frames, annotated from scratch via multi-round expert labeling) to serve as a high-confidence benchmark.
Figure 3: G1 Domain-Specific Hybrid Dataset Construction Pipeline. Our approach synergizes cost-free, high-fidelity 3D/2D geometry rendered from a 9-camera simulation environment (Isaac Lab) with Human-in-the-Loop (HITL) annotations to correct visual domain shifts, enabling robust native pose estimation.